### v1

Abstract
This paper addresses distributed optimization problems in non-i.i.d. scenarios, focusing on the interplay between communication and computation efficiency. To this end, we propose FlexGT, a flexible snapshot gradient tracking method with tunable numbers of local updates and neighboring communications in each round. Leveraging a unified convergence analysis framework, we prove that FlexGT achieves a linear or sublinear convergence rate depending on objective-specific properties–from (strongly) convex to nonconvex–and the above-mentioned tunable parameters. FlexGT is provably robust to the heterogeneity across nodes and attains the best-known communication and computation complexity among existing results. Moreover, we introduce an accelerated gossip-based variant, termed Acc-FlexGT, and show that with prior knowledge of the graph, it achieves a Pareto-optimal trade-off between communication and computation. Particularly, Acc-FlexGT achieves the optimal iteration complexity of $\tilde{\mathcal{O}}\left(L/\epsilon+L\sigma^{2}/\left(n\epsilon^{2}\sqrt{1-\sqrt{\rho_{W}}}\right)\right)$ for the nonconvex case, matching the existing lower bound up to a logarithmic factor, and improves the existing results for the strongly convex case by a factor of $\tilde{\mathcal{O}}\left(1/\sqrt{\epsilon}\right)$, where $\epsilon$ is the targeted accuracy, $n$ the number of nodes, $L$ the Lipschitz constant, $\rho_{W}$ the spectrum gap of the graph, and $\sigma$ the stochastic gradient variance. Numerical examples are provided to demonstrate the effectiveness of the proposed methods.
keywords:
Distributed optimization, data heterogeneity, complexity analysis, gradient tracking, accelerated communication, Pareto-optimal
{MSCcodes}
90C06, 90C25, 90C26
1 Introduction
With rapid growth in both data volume and device scale [21], distributed optimization methods have gained increasing attention due to their wide range of applications such as in cooperative control [33], distributed sensing [1], large-scale machine learning [49], distributed training of large language models (LLMs) [16], to name a few. In this paper, we consider a distributed stochastic optimization problem jointly solved by $n$ nodes over a network:
(1) $\underset{x\in\mathbb{R}^{p}}{\min}f\left(x\right)=\frac{1}{n}\sum_{i=1}^{n}{\underset{:=f_{i}\left(x\right)}{\underbrace{\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}\left[f_{i}\left(x;\xi_{i}\right)\right]}}},$
where $x\in\mathbb{R}^{p}$ is the global decision variable. The objective function $f:\mathbb{R}^{p}\rightarrow\mathbb{R}$ is the sum of $n$ local objective functions $f_{i}$, $i=1,\dots,n$, conditioned on the local data sample $\xi_{i}$ with distribution $\mathcal{D}_{i}$.
1.1 Literature Review
Decentralized stochastic gradient descent (DSGD) [39] is a simple yet efficient algorithm for solving problem (1). It performs stochastic gradient descent locally while communicating with neighbors to achieve consensus, which enables data parallelism and avoids communication bottlenecks in centralized architectures [32]. However, the native DSGD methods often struggle in scenarios with non-independent and identically distributed (non-i.i.d.) local datasets [24, 12]. For (1), such a scenario corresponds to that the distribution $\mathcal{D}_{i}$ is different. This data heterogeneity induces gradient drift in the sense that local gradients computed at each node deviate from the global gradient direction, necessitating decaying stepsizes to ensure exact convergence [35, 24]. This requirement limits the applicability of DSGD in many real-world tasks, as non-i.i.d. data distributions are common.
Many efforts have been made to address this issue of data heterogeneity in distributed optimization. Gradient tracking (GT) methods have gained popularity for mitigating the effect [51, 11, 41, 23, 46, 48]. Notably, Xu et al. [51] proposed AugDGM, which introduces an auxiliary variable at each node to track the average of the local gradients, enabling exact convergence with a constant stepsize. This method has been extended to broader settings, including time-varying graphs [34] and directed graphs [38, 40], with convergence guarantees. Pu et al. [37] extended GT to the stochastic setting, introducing distributed stochastic gradient tracking (DSGT). They proved that DSGT achieves linear convergence to a neighborhood of the optimal solution, determined by the stochastic gradient noise. This extension enhances the applicability of GT, for instance, to many machine learning tasks. Although GT methods are robust to data heterogeneity, they exhibit a stronger dependence on the graph compared to DSGD [13, 18], leading to increased communication overhead. The consensus alternating direction method of multipliers (C-ADMM) [27] and its inexact variant [6] also exhibit robustness to data heterogeneity. However, they require either solving a subproblem exactly at each iteration, which incurs high computation overhead, or assuming a well-structured objective function to enable linear approximations, which limits their applicability. These observations highlight the need for more efficient algorithms, especially in resource-constrained scenarios that involve both communication and computation.
To mitigate the high communication burden, many improved GT-based methods have been developed [36, 30, 43, 10, 42]. A popular strategy is to skip some communication steps and perform local updates only. For instance, FedAvg [32] (also known as Local SGD [44]) is widely used in federated learning [8, 17], where periodic communication occurs between server and worker nodes, with partial node participation to reduce communication overhead. In decentralized settings, Nguyen et al. [36] introduced LU-GT, a variant of DSGT that uses multiple local updates, and provided communication complexity results matching those of Local SGD for nonconvex objectives. Subsequently, Liu et al. [30] proposed the gradient-sum-tracking method K-GT, which reduces communication complexity by achieving lower stochastic gradient variance. However, both [36] and [30] overlook the side effect of how the local updates increase computation complexity due to worse inter-node consensus. Another approach is to randomly or periodically communicate over a subset of the graph’s edges to reduce communication [39, 25, 26, 52]. These methods can be seen as approximate asynchronous update schemes, which often require specific connectivity assumptions [39] or particular topologies, such as exponential graphs [52].
In contrast to directly skipping certain communications as in Local SGD [44], another strategy involves using accelerated consensus protocols. For instance, Chebyshev acceleration [43, 10] and accelerated gossip [28, 31, 54] methods enable the distributed algorithm to behave more like a centralized algorithm. In particular, Song et al. [43] introduced an optimal gradient tracking (OGT) method with Chebyshev acceleration for deterministic optimization problems with smooth and strongly convex objective functions. Leveraging prior knowledge of the network topology, they demonstrated that a limited number of communication steps per round could enable the distributed algorithm to replicate centralized updates. Building on this, the stochastic variant SS-DSGT [10] was proposed, achieving lower communication complexity compared to DSGT. For nonconvex cases, Lu et al. [31] developed an accelerated DSGT method called DeTAG, achieving optimal iteration complexity. However, the effects of these acceleration techniques on both communication and computation remain unclear. Although communication quantization and compression can be used to further reduce the amount of data transmitted in each round [20, 42], their impacts on algorithmic accuracy must be carefully considered. For a comprehensive overview of other efforts to improve communication efficiency, readers are referred to recent survey papers [47, 5].
Some recent efforts have focused on both the communication and computation efficiency of distributed optimization algorithms. Berahas et al. [3] proposed a distributed gradient descent method with multiple communication steps in each round, named NEAR-DGD, and evaluated the performance of the algorithm using a weighted combination of the number of communication and computation steps. In [2], the authors explored the use of multiple local updates and inter-node communication in each round with a deterministic GT method, showing that increasing the number of communication steps improves the convergence rate in strongly convex settings. However, in distributed stochastic optimization, balancing communication and computation is more challenging. This is because stochastic gradient noise significantly influences the convergence rate, making the results from deterministic methods [3, 2] inapplicable to stochastic settings. Although the results in [15, 29] highlight the impact of communication and computation protocols on iteration complexity, they do not provide a specific trade-off strategy between these two factors and are all affected by the presence of data heterogeneity. Additionally, most of the studies focus on specific types of objective functions, such as strongly convex or nonconvex objectives, or rely on strict assumptions, such as bounded gradients [29]. We summarize and compare the most relevant algorithms in Table 1.
Part of the results in this work were previously presented at a conference [14]. The current work: i) extends the algorithm in [14] by incorporating snapshot GT and accelerated gossip communication, resulting in a more flexible framework with improved convergence guarantees; ii) develops a unified theoretical framework for strongly convex, convex, and nonconvex objective functions with comprehensive complexity results; and iii) includes more extensive numerical experiments on both synthetic and real-world datasets.
Table 1: Relevant algorithms for solving Problem (1)
Ref. Objectives Comm. Comp. Stoch. Acc. Non-i.i.d.
DSGT[37, 18] SC, Co, NC Single Single
SS-GT [43] SC Single Single
SS-DSGT [10] SC Single Single
LU-GT[36] NC Single Multi.
K-GT [30] NC Single Multi.
DeTAG [31] NC Multi. Single
GTA [2] SC Multi. Multi.
S-NEAR-DGD[15] SC Multi. Multi.
DFL [29] NC Multi. Multi.
(Acc-)FlexGT (ours) SC, Co, NC Multi. Multi.
  • Note: SC, Co, and NC denote strongly convex, convex, and nonconvex objective functions, respectively. “Single” and “Multi.” indicate that the algorithm uses single or multiple local communication (Comm.) or computation (Comp.) steps per round, respectively. “Stoch.” refers to the use of stochastic gradients. “Acc.” refers to communication acceleration.
1.2 Contributions
This paper solves the distributed stochastic optimization problem (1) in scenarios with non-i.i.d local datasets. We introduce the FlexGT algorithm and illustrate its trade-offs between communication and computation. Pareto-optimal trade-offs are further achieved through an accelerated variant, termed Acc-FlexGT. The main contributions are summarized as follows:
  • We propose FlexGT, a flexible gradient tracking method that allows adjustable numbers of local updates and communications in each round, enabling trade-offs between communication and computation. By introducing a so-called snapshot point within the GT step, FlexGT mitigates the inconsistency between GT variables during local updates, resulting in improved convergence. Moreover, by integrating an accelerated gossip communication protocol, we develop Acc-FlexGT, which significantly enhances both communication and computation efficiency, and, to the best of our knowledge, is the first to achieve Pareto-optimal trade-offs between these two aspects, under the assumption of prior knowledge of the graph.
  • We develop a unified convergence analysis framework for FlexGT and Acc-FlexGT, covering objective functions that are strongly convex, convex, or nonconvex. We establish linear or sublinear convergence rates with the best-known communication and computation complexity in most settings, in terms of their dependencies on problem-specific parameters, compared to existing results (Theorem 1 and Corollary 1). In particular, we prove that Acc-FlexGT achieves the optimal iteration complexity, matching the existing lower bound in the nonconvex case up to a logarithmic factor, and improves existing results in the strongly convex case (Corollary 2). To the best of our knowledge, this work provides the first complexity analysis of an accelerated GT method in a convex setting. This unified analysis framework can be extended to a broader class of distributed optimization algorithms, which is of independent interest.
1.3 Paper Organization
The remainder of this paper is structured as follows. Section 2 formulates the distributed stochastic optimization problem over networks under standard assumptions and introduces the FlexGT and Acc-FlexGT algorithms. Section 3 then presents the main convergence result of the algorithms for strongly convex, convex, and nonconvex objective functions, showing the dependency on both the amount of communication and computation. Section 4 provides the proof of the result. Numerical experiments on synthetic and real-world datasets are given in Section 5 to validate the effectiveness of FlexGT and Acc-FlexGT. Finally, Section 6 concludes the paper.
1.4 Notations
[…]
2 Problem Formulation and Algorithm Design
To solve the distributed stochastic optimization problem (1) in a decentralized manner, we consider the following equivalent problem:
(2) $\underset{\mathbf{x}\in\mathbb{R}^{n\times p}}{\min}F\left(\mathbf{x}\right)=\frac{1}{n}\sum_{i=1}^{n}{\underset{:=f_{i}\left(x_{i}\right)}{\underbrace{\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}\left[f_{i}\left(x_{i};\xi_{i}\right)\right]}},}\quad\mathrm{s.t.}\,\,x_{i}=x_{j},\quad\forall i,j\in\left[n\right]$
where $\mathbf{x}:=\left[x_{1},x_{2},\dots,x_{n}\right]\in\mathbb{R}^{n\times p}$ denotes the collection of the local copies of the decision variables from all nodes. The constraints guarantee that problems (1) and (2) are equivalent. Moreover, the agents are connected over a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Here, $\mathcal{V}=\{1,2,...,n\}$ represents the set of nodes, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the set of edges consisting of ordered pairs $(i,j)$ representing the communication link from node $j$ to node $i$. For node $i\in[n]$, we define $\mathcal{N}_{i}=\{j|\left(i,j\right)\in\mathcal{E}\}$ as the set of its neighboring nodes (including $i$ itself).
2.1 Assumptions
We make the following assumptions regarding the objective function $f_{i}$, its gradient $\nabla f_{i}$, and the graph $\mathcal{G}$.
Assumption 1 (Convexity).
For each $f_{i}$, there is a constant $\mu\geqslant 0$ such that for all $x,x^{\prime}\in\mathbb{R}^{p}$,
$f_{i}\left(x^{\prime}\right)-f_{i}\left(x\right)\geqslant\left<\nabla f\left(x\right),x^{\prime}-x\right>+\frac{\mu}{2}\left\|x^{\prime}-x\right\|^{2}.$
Particularly, we say that $f_{i}$ is $\mu$-strongly convex if $\mu>0$; otherwise, we refer to it as convex.
Assumption 2 (Smoothness).
Each $f_{i}$ is $L$-smooth, i.e., there is a Lipschitz constant $L>0$ such that for all $x,x^{\prime}\in\mathbb{R}^{p}$,
$f_{i}\left(x^{\prime}\right)-f_{i}\left(x\right)\leqslant\left<\nabla f\left(x\right),x^{\prime}-x\right>+\frac{L}{2}\left\|x-x^{\prime}\right\|^{2}.$
In our algorithms, we assume that each agent $i$ obtains an unbiased noisy gradient of the form $\nabla f_{i}\left(x;\xi_{i}\right)$ by querying a stochastic oracle ($\mathcal{S}\mathcal{O}$) with $\xi_{i}\sim\mathcal{D}_{i}$ at each iteration, where ${D}_{i}$ is an arbitrary distribution.
Assumption 3 (Bounded variance).
For the stochastic gradient $\nabla f_{i}\left(x;\xi_{i}\right)$ with $\xi_{i}\sim\mathcal{D}_{i}$, there exist a constant $\sigma\geqslant 0$ such that for all $x\in\mathbb{R}^{p}$,
$\mathbb{E}\left[\left\|\nabla f_{i}\left(x;\xi_{i}\right)-\nabla f_{i}\left(x\right)\right\|^{2}\right]\leqslant\sigma^{2}.$
Assumption 4 (Graph connectivity).
The weight matrix $W$ induced by the graph $\mathcal{G}$ is doubly stochastic, i.e., $W\mathbf{1}=\mathbf{1},\mathbf{1}^{\top}W=\mathbf{1}^{\top}$, and $\rho_{W}:=\left\|W-\mathbf{J}\right\|_{2}^{2}<1.$
Note that a suitable weight matrix $W$ satisfying Assumption 4 can be constructed for any undirected and connected graph. For example, a weight matrix can be readily determined using the Metropolis-Hastings protocol [50].
2.2 FlexGT and Acc-FlexGT Algorithms
Now, we proceed to present the FlexGT algorithm and its accelerated variant, Acc-FlexGT, for solving problem (2). The pseudo-code for the algorithms is provided in Algorithm 1. Note that the integers $\alpha,\beta\geqslant 1$ denote the number of communication and computation steps performed in each round, respectively. A flow diagram illustrating the communication and computation of FlexGT is shown in Fig. 1.
Algorithm 1 FlexGT and Acc-FlexGT
0: Initial points $z_{i,0}=x_{i,0}\in\mathbb{R}^{p}$ and $y_{i,0}=\nabla_{x}f_{i}\left(x_{i,0};\xi_{i,0}\right)$, number of communication and computation steps $\alpha,\beta\geqslant 1$, and stepsize $\gamma>0$.
1:for round $k=0,1,\dots$, each node $i\in[n]$do
2:  Update snapshot point: $z_{i,\beta k}=x_{i,\beta k}$.
3:  for $l=\beta k,\beta k+1,\dots,\beta\left(k+1\right)-1$ do
4:   Obtain an unbiased noisy gradient $\nabla f_{i}\left(z_{i,l+1};\xi_{i,l+1}\right)$ by querying $\mathcal{S}\mathcal{O}$.
5:   Perform local updates:
$x_{i,l+1}$ $=x_{i,l}-\gamma y_{i,l}\quad\quad\text{\# gradient descent;}$
$z_{i,l+1}$ $=z_{i,l}\quad\quad\text{\# snapshot point;}$
$y_{i,l+1}$ $=y_{i,l}+\nabla f_{i}\left(z_{i,l+1};\xi_{i,l+1}\right)-\nabla f_{i}\left(z_{i,l};\xi_{i,l}\right)\quad\text{\# gradient tracking}$
6:  end for
7:  Let $\theta_{i,0}=\theta_{i,-1}=\left\{x_{i,\beta\left(k+1\right)},y_{i,\beta\left(k+1\right)}\right\}$.
8:  for $s=0,1,\dots,\alpha-1$ do
9:   FlexGT: Direct inter-node communication:
$\vskip-8.5359pt\theta_{i,s+1}=\sum_{j\in\mathcal{N}_{i}}{w_{ij}\theta_{j,s}}.$
10:   Acc-FlexGT: Accelerated gossip communication:
$\vskip-2.84544pt\theta_{i,s+1}=\left(1+\eta\right)\sum_{j\in\mathcal{N}_{i}}{w_{ij}\theta_{j,s}-\eta}\theta_{i,s-1},$
         where $\eta=\frac{1-\sqrt{1-\rho_{W}}}{1+\sqrt{1-\rho_{W}}}$.
11:  end for
12:end for
Figure 1: Flow diagram illustrating the computation and communication of FlexGT. The variable $\psi_{i,l}$, $i=1,\dots,n$, represents the collection of $x_{i,l}$, $y_{i,l}$, and $z_{i,l}$. Only $x_{i,l}$ and $y_{i,l}$ are communicated. The communication is weighted by the matrix $W$.
We introduce the following notations to represent the collection of decision parameters, GT variables, and snapshot points, respectively:
$\mathbf{x}_{t}=\left[x_{1,t},x_{2,t},\dots,x_{n,t}\right]^{\top}\in\mathbb{R}^{n\times p},$
$\mathbf{y}_{t}=\left[y_{1,t},y_{2,t},\dots,y_{n,t}\right]^{\top}\in\mathbb{R}^{n\times p},$
$\mathbf{z}_{t}=\left[z_{1,t},z_{2,t},\dots,z_{n,t}\right]^{\top}\in\mathbb{R}^{n\times p},$
Then, Algorithm 1 can be rewritten in a compact form:
(3a) $\mathbf{x}_{\beta\left(k+1\right)}$ $=\bar{W}\left(\mathbf{x}_{\beta k}-\gamma\sum_{j=0}^{\beta-1}{\mathbf{y}_{\beta k+j}}\right),$
(3b) $\mathbf{y}_{\beta\left(k+1\right)}$ $=\bar{W}\left(\mathbf{y}_{\beta k}+\nabla G_{\beta\left(k+1\right)}-\nabla G_{\beta k}\right),$
where
$\nabla G_{t}:=\left[\dots,\nabla f_{i}\left(z_{i,t};\xi_{i,t}\right),\dots\right]^{\top}\in\mathbb{R}^{n\times p}$
denotes the collection of the stochastic gradient at the snapshot point $\mathbf{z}_{t}$ at iteration $t$, and its expectation is denoted by
$\nabla F_{t}:=\mathbb{E}\left[\nabla G_{t}\right]=\left[\dots,\nabla f_{i}\left(z_{i,t}\right),\dots\right]^{\top}\in\mathbb{R}^{n\times p}.$
We note that $\bar{W}$ takes different forms depending on the communication protocol employed. In the case of FlexGT, we have $\bar{W}=W^{\alpha}$. Otherwise, when the accelerated gossip communication protocol is used as in Acc-FlexGT, we have $\bar{W}=M_{\alpha}$ with $M_{\alpha}$ obtained through $\alpha$ communication steps:
$M_{s+1}=\left(1+\eta\right)WM_{s}-\eta M_{s-1},\quad s\in\{0,1,\dots,\alpha-1\}$
with $M_{-1}=M_{0}=I$. Although in both cases there are $\alpha$ communications per round, the difference is in prior knowledge, e.g., the spectral gap of the graph, which leads to distinct communication properties, as described in the following lemma.
Lemma 1.
Suppose Assumption 4 holds and denote $\bar{\rho}_{W}:=\left\|\bar{W}-\mathbf{J}\right\|^{2}$. Then, for FlexGT, we have
(4) $\bar{\rho}_{W}=\left\|W^{\alpha}-\mathbf{J}\right\|^{2}\leqslant\rho_{W}^{\alpha};$
and for Acc-FlexGT,
(5) $\bar{\rho}_{W}=\left\|M_{\alpha}-\mathbf{J}\right\|^{2}\leqslant 2\left(1-\sqrt{1-\sqrt{\rho_{W}}}\right)^{2\alpha}.$
Equation (4) is straightforward to derive based on Assumption 4, and (5) was stated at Proposition 3 in [28].
Regarding the snapshot point $\mathbf{z}$, we have
(6) ${\mathbf{z}_{\beta k+t}}=\,\,\mathbf{x}_{\beta k},\quad t=0,1,\dots,\beta-1,$
which implies that $\mathbf{y}$ tracks the stochastic gradient at the snapshot point $\mathbf{z}$ instead of the current decision variable $\mathbf{x}$, i.e.,
(7) $\mathbf{1}^{\top}{\mathbf{y}_{\beta k+t}}=\mathbf{1}^{\top}{\nabla G_{\beta k+t}}=\sum_{i=1}^{n}{\nabla f_{i}\left(x_{i,\beta k};\xi_{i,\beta k+t}\right)}.$
This prevents the tracking variables from diverging excessively due to multiple local updates, i.e., the gradient difference $\nabla G_{t+1}-\nabla G_{t}$ remains controlled, which is crucial for stability in the presence of heterogeneous data. This method differs from the existing snapshot GT approach [43, 10], which randomly skips updating the tracking variables while performing their consensus at each iteration, potentially leaving computation resources idle. In contrast, FlexGT utilizes the stochastic gradient at each iteration to slightly update the tracking variables and perform communication periodically to avoid extra costs from redundant initialization.
The FlexGT algorithm generalizes many existing methods by incorporating an adjustable communication and computation protocol parameterized by $\alpha$ and $\beta$. Specifically, it recovers DSGT [37] ($\alpha=\beta=1$), LU-GT [36] ($\alpha=1,\beta\geqslant 1$), and DeTAG ($\alpha\geqslant 1,\beta=1$) as special cases. It can also recover DFL [29] when an independent weight matrix is used for updating $\mathbf{y}$. This highlights the flexibility of FlexGT and its applicability to a broader range of problems.
3 Main Results
In this section, we present the main convergence results of the FlexGT and Acc-FlexGT algorithms for strongly convex, convex, and nonconvex objective functions under a unified analysis framework. To this end, we first define the following Lyapunov function for (strongly) convex cases, consisting of the optimality gap, consensus error, and GT error:
(8) $V_{t}=\left\|\bar{x}_{t}-x^{*}\right\|^{2}+c_{x}\left\|\tilde{\mathbf{x}}_{t}\right\|^{2}+c_{y}\left\|\tilde{\mathbf{y}}_{t}\right\|^{2},$
where $x^{*}$ indicates the global optimal solution of problem (1), $t$ denotes the iteration, $c_{x}$ and $c_{y}$ are coefficients to be properly designed later, and
$\tilde{\mathbf{x}}_{t}=\mathbf{x}_{t}-\mathbf{1}\bar{x}_{t},\quad\bar{x}_{t}=\mathbf{1}^{\top}\mathbf{x}_{t}/n,\quad\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}-\mathbf{1}\bar{y}_{t},\quad\bar{y}_{t}=\mathbf{1}^{\top}\mathbf{y}_{t}/n.$
Table 2: Communication and computation complexity of FlexGT for strongly convex, convex, and nonconvex objective functions.
[…]
SC $\frac{\beta L}{\left(1-\rho_{W}^{\alpha}\right)^{2}\mu}+\frac{\sigma^{2}}{\mu^{2}n\epsilon}+\frac{\beta\sqrt{L\rho_{W}^{\alpha}\sigma^{2}}}{\mu^{3/2}\left(1-\rho_{W}^{\alpha}\right)^{3/2}\epsilon^{1/2}}$ $\frac{\alpha L}{\left(1-\rho_{W}^{\alpha}\right)^{2}\mu}+\frac{\alpha\sigma^{2}}{\beta\mu^{2}n\epsilon}+\frac{\alpha\sqrt{L\rho_{W}^{\alpha}\sigma^{2}}}{\mu^{3/2}\left(1-\rho_{W}^{\alpha}\right)^{3/2}\epsilon^{1/2}}$
FlexGT Co $\frac{\beta L}{\left(1-\rho_{W}^{\alpha}\right)^{2}\epsilon}+\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\beta\sqrt{L\rho_{W}^{\alpha}\sigma^{2}}}{\left(1-\rho_{W}^{\alpha}\right)^{3/2}\epsilon^{3/2}}$ $\frac{\alpha L}{\left(1-\rho_{W}^{\alpha}\right)^{2}\epsilon}+\frac{\alpha\sigma^{2}}{n\beta\epsilon^{2}}+\frac{\alpha\sqrt{L\rho_{W}^{\alpha}\sigma^{2}}}{\left(1-\rho_{W}^{\alpha}\right)^{3/2}\epsilon^{3/2}}$
NC $\frac{L\sigma^{2}}{n\epsilon^{2}}+\frac{\beta L}{\left(1-\rho_{W}^{\alpha}\right)^{2}\epsilon}+\frac{\beta R_{0}}{\epsilon}+\frac{\beta L\sqrt{\rho_{W}^{\alpha}\sigma^{2}}}{\left(1-\rho_{W}^{\alpha}\right)^{3/2}\epsilon^{3/2}}$ $\frac{\alpha L\sigma^{2}}{n\beta\epsilon^{2}}+\frac{\alpha L}{\left(1-\rho_{W}^{\alpha}\right)^{2}\epsilon}+\frac{\alpha R_{0}}{\epsilon}+\frac{\alpha L\sqrt{\rho_{W}^{\alpha}\sigma^{2}}}{\left(1-\rho_{W}^{\alpha}\right)^{3/2}\epsilon^{3/2}}$
Table 3: Communication and computation complexity of Acc-FlexGT for strongly convex, convex, and nonconvex objective functions. The last column presents the corresponding best-known or optimal iteration complexity.
[…]
SC $\beta\frac{L}{\mu}+\frac{\sigma^{2}}{n\mu^{2}\epsilon}$ $\left(\frac{\sigma^{2}}{n\beta\mu^{2}\epsilon}+\frac{L}{\mu}\right)\frac{1}{\sqrt{1-\sqrt{\rho_{W}}}}$ $\frac{\sigma^{2}}{n\mu^{2}\epsilon}+\frac{\sqrt{L\sigma^{2}}}{\mu\sqrt{1-\sqrt{\rho_{W}}}\epsilon^{1/2}}$ [10] a
Acc- Co $\frac{L\sigma^{2}}{n\epsilon^{2}}+\frac{\beta\sqrt{L}}{\epsilon}$ $\left(\frac{L\sigma^{2}}{n\beta\epsilon^{2}}+\frac{\sqrt{L}}{\epsilon}\right)\frac{1}{\sqrt{1-\sqrt{\rho_{W}}}}$ / b
FlexGT
NC $\frac{L\sigma^{2}}{n\epsilon^{2}}+\frac{\beta L}{\epsilon}$ $\left(\frac{L\sigma^{2}}{n\beta\epsilon^{2}}+\frac{L}{\epsilon}\right)\frac{1}{\sqrt{1-\sqrt{\rho_{W}}}}$ $\frac{L\sigma^{2}}{n\epsilon^{2}}+\frac{L}{\sqrt{1-\sqrt{\rho_{W}}}\epsilon}$ [31] c
  • a
    To the best of our knowledge, this is the best-known result in the existing literature in this setting.
  • b
    The mark "/" indicates no relevant result in the existing literature.
  • c
    This is the optimal iteration complexity that matches the existing lower bound up to a logarithmic factor.
3.1 Convergence
The following theorem presents the convergence results of FlexGT and Acc-FlexGT for strongly convex, convex, and nonconvex objective functions under specific assumptions.
Theorem 1.
Suppose Assumptions 24 hold. There exists a constant stepsize
(9) $\gamma=\mathcal{O}\left(\min\left\{1,\frac{\left(1-\bar{\rho}_{W}\right)^{2}}{\bar{\rho}_{W}}\right\}\right)\frac{1}{\beta L}$
such that the optimization error is bounded as follows:
i) Strongly convex: Under the additional Assumption 1 with $\mu>0$, we have for all $k\geqslant 0$
(10) $\mathbb{E}\left[V_{\beta\left(k+1\right)}\right]$ $\leqslant\underset{\mathrm{linear\,\,rate}}{\underbrace{\left(1-\min\left\{\frac{\mu\beta\gamma}{2},\frac{1-\bar{\rho}_{W}}{8}\right\}\right)}}\mathbb{E}\left[V_{\beta k}\right]$
$\quad+\underset{\mathrm{centralized\,\,error}}{\underbrace{\frac{\gamma^{2}\beta\sigma^{2}}{n}}}+\underset{\mathrm{topology\,\,effect}}{\underbrace{\frac{1664\gamma^{3}\beta^{3}L\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}}}.$
ii) Convex: Under the additional Assumption 1 with $\mu\geqslant 0$, we have for all $K\geqslant 1$,
(11) $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right]}$ $\leqslant\underset{\mathrm{sublinear\,\,rate}}{\underbrace{\frac{2\mathbb{E}\left[V_{0}\right]}{\gamma\beta K}}}+\underset{\mathrm{centralized\,\,error}\mathrm{}}{\underbrace{\frac{2\gamma\sigma^{2}}{n}}}$
$\quad+\underset{\mathrm{topology\,\,effect}}{\underbrace{\frac{3328\gamma^{2}\beta^{2}L\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}}}.$
iii) Nonconvex: It holds that for all $K\geqslant 1$,
(12) $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]}$ $\leqslant\underset{\mathrm{sublinear\,\,rate}}{\underbrace{\frac{8\mathbb{E}\left[f\left(\bar{x}_{0}\right)-f^{*}\right]}{\gamma\beta K}+\frac{8\gamma^{2}\beta^{2}L^{2}\bar{\rho}_{W}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]}{n\left(1-\bar{\rho}_{W}\right)^{3}K}}}$
$\quad+\underset{\mathrm{centralized\,\,error}}{\underbrace{\frac{4\gamma L\sigma^{2}}{n}}}+\underset{\mathrm{topology\,\,effect}}{\underbrace{\frac{3328\gamma^{2}\beta^{2}L^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}}}.$
Proof.
See Section 4.
Theorem 1 shows that for the strongly convex case, with constant stepsize $\gamma$, FlexGT converges linearly to a neighborhood of the optimal solution of problem (1). The steady-state error consists of two parts: one is the term of stochastic gradient noise matching that of the centralized SGD algorithm, while the other part is due to the decentralized graph topology. Similarly, in the convex case with $\mu=0$, the rate is sublinear in the sense of the running sum of the objective function. Finally, in the nonconvex case, the running sum of the gradient norm of FlexGT converges to a neighborhood of a stationary point at a sublinear rate. Note that these results highlight a clear dependence on the parameters $\mu$ and $L$ of the objective function, the computation steps $\beta$, and the spectrum gap $\bar{\rho}_{W}$ related to communication steps $\alpha$ and $\rho_{W}$ in Assumption 4.
Remark 1.
Compared to the existing results in [29, 15], FlexGT handles data heterogeneity [25], i.e. $\forall x$,
$\zeta_{f}:=\sup\left\{\frac{1}{n}\sum_{i=1}^{n}\left\|\nabla f_{i}(x)-\nabla f(x)\right\|^{2}\right\},$
and does not require the assumption of uniformly bounded stochastic gradients [29], making it more robust in non-i.i.d. settings. Compared to the best-known results for DSGT ($\alpha=\beta=1$) without acceleration in [18], our findings recover their results in all three settings  [^1] Under Assumption 4 on the graph and using the same convergence metric as in this work. and extend to more general cases with $\alpha\geqslant 1$ and $\beta\geqslant 1$. Note that K-GT [30] is a special case of FlexGT with $\alpha=1$ and $\beta>1$, and the communication complexity obtained in this work improves upon that of K-GT by a factor of $1/\sqrt{1-\rho_{W}}$ in terms of the topology effect for the nonconvex setting, while K-GT experiences lower stochastic gradient noise.
Based on the convergence results in Theorem 1, we can further derive the iteration complexity of FlexGT for different objective functions as follows.
Corollary 1.
Suppose Assumptions 24 hold. Let the stepsize satisfy the condition (9) given in Theorem 1, and denote by $\hat{K}$ the number of iterations required for FlexGT to reach an arbitrary accuracy $\epsilon>0$.
i) Strongly convex: Under the additional Assumption 1 with $\mu>0$, we have
(13) $\hat{K}={\mathcal{O}}\left(\frac{L}{\left(1-\bar{\rho}_{W}\right)^{2}\mu}\log\frac{\left[V_{0}\right]}{\epsilon}+\frac{\sigma^{2}}{\beta\mu^{2}n\epsilon}+\frac{\sqrt{L\bar{\rho}_{W}\sigma^{2}}}{\mu^{3/2}\left(1-\bar{\rho}_{W}\right)^{3/2}\epsilon^{1/2}}\right).$
ii) Convex: Under the additional Assumption 1 with $\mu\geqslant 0$, we have
(14) $\hat{K}=\mathcal{O}\left(\frac{L}{\left(1-\bar{\rho}_{W}\right)^{2}\epsilon}+\frac{\sigma^{2}}{\beta n\epsilon^{2}}+\frac{\sqrt{L\bar{\rho}_{W}\sigma^{2}}}{\left(1-\bar{\rho}_{W}\right)^{3/2}\epsilon^{3/2}}\right)\mathbb{E}\left[V_{0}\right].$
iii) Nonconvex: It holds
(15) $\hat{K}=\mathcal{O}$ $\left(\frac{L}{\left(1-\bar{\rho}_{W}\right)^{2}\epsilon}+\frac{R_{0}}{\epsilon}+\frac{L\sigma^{2}}{\beta n\epsilon^{2}}\right.\left.+\frac{L\sqrt{\bar{\rho}_{W}\sigma^{2}}}{\left(1-\bar{\rho}_{W}\right)^{3/2}\epsilon^{3/2}}\right)\left(\mathbb{E}\left[f\left(\bar{x}_{0}\right)-f^{*}\right]\right),$
where $R_{0}:=\frac{1}{n}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]\left(\mathbb{E}\left[f\left(\bar{x}_{0}\right)-f^{*}\right]\right)^{-1}$.
Proof.
Due to space constraints, we present only the main ideas of the proof while omitting the detailed steps. Denote by $\hat{\gamma}$ the obtained upper bound of the stepsize in (9) and
$\varUpsilon_{1}:=\frac{\beta\sigma^{2}}{n\upsilon_{0}},\quad\varUpsilon_{2}:=\frac{\beta^{3}L\bar{\rho}_{W}\sigma^{2}}{\upsilon_{0}\left(1-\bar{\rho}_{W}\right)^{3}},$
the constant part of the centralized error and the topology effect part in (10)–(12) controlled by the stepsize $\gamma$, respectively, where $\upsilon_{0}$ denotes either $\mathbb{E}\left[V_{0}\right]$ or $\mathbb{E}\left[f\left(\bar{x}_{0}\right)-f^{*}\right]$, depending on the specific case under consideration. Then, we choose a sufficiently small $\gamma$ such that the centralized error and topology effect part have the same order as the linear or sublinear part in terms of the number of iterations $\hat{K}$. To this end, letting the stepsize $\gamma$ satisfy
$\min\left\{\frac{\ln\left(\max\left\{2,\beta^{2}\mu^{2}\hat{K}/\varUpsilon_{1},\beta^{3}\mu^{3}\hat{K}^{2}/\varUpsilon_{2}\right\}\right)}{\beta\mu\hat{K}},\,\,\hat{\gamma}\right\}$
for the strongly convex case; and
$\min\left\{\frac{1}{\sqrt{\varUpsilon_{1}\hat{K}}},\,\,\frac{1}{\sqrt[3]{\varUpsilon_{2}\hat{K}}},\,\,\hat{\gamma}\right\}$
for convex and nonconvex cases, we obtain the complexity results. Similar techniques can be found in [45, 19, 13].
Based on the iteration complexity derived in Corollary 1 and the number of local updates and communications per iteration, we can derive the communication and computation complexity of FlexGT for different objective functions, as summarized in Table 2. To the best of our knowledge, FlexGT achieves the best-known complexity among GT-based methods for solving problem in most settings (cf. Remark 1).
Remark 2.
Table 2 highlights the impact of $\alpha$ and $\beta$ on the communication and computation complexity of FlexGT. Unlike previous work, our results reveal an inherent trade-off between communication and computation. Particularly, as the number of computation steps $\beta$ increases, the communication complexity decreases while the computation complexity increases. Conversely, as the number of communication steps $\alpha$ increases, the computation complexity decreases, while the communication complexity may either increase or decrease depending on $\rho_{W}$. This highlights the need for a careful design of $\alpha$ and $\beta$ to balance communication and computation costs according to the specific application requirements.
3.2 Pareto-optimal Trade-offs
Figure 2: The number of communication and computation steps needed for FlexGT to achieve an accuracy of $\epsilon=10^{-4}$ with $\alpha,\beta=1,2,\dots,100$. Pareto-optimal solutions (red lines) are achieved with $\alpha=4$ and $\alpha=32$ for strongly convex (left) and nonconvex (right) cases, respectively.
The trade-off between communication and computation is illustrated in Fig. 2, where each gray point reflects a specific choice of $\alpha$ and $\beta$ in the range from 1 to 100. The coordinates of each point represent the estimated number of communication steps (horizontal axis) and number of computation steps (vertical axis) required to achieve an accuracy of $\epsilon$, obtained from Table 2. It can be observed that, according to the definition of a Pareto-optimal solution in multi-objective optimization, a specific set of parameters lies on the Pareto frontier, as shown by the bold red lines. This set achieves the lowest overall complexity for reaching a given accuracy level, meaning that improvements in either communication or computation complexity are not possible without compromising the other.
Definition 1 (Pareto-optimality [4, 7]).
For a multi-objective optimization problem that minimizes $\left\{g_{i}\left(\omega\right)\right\}_{j\in\left[m\right]}$, a point $\omega$ is Pareto-optimal if there does not exist another point $\omega^{\prime}$ such that: $\forall j\in[m]$, $g_{j}\left(\omega^{\prime}\right)\leqslant g_{j}\left(\omega\right)$; and $\exists j\in\left[m\right]$, $g_{j}\left(\omega^{\prime}\right)<g_{j}\left(\omega\right)$. The set of Pareto-optimal points is called the Pareto frontier.
The following corollary shows that by appropriately incorporating accelerated gossip communication, Acc-FlexGT substantially improves the iteration complexity and achieves Pareto-optimal trade-offs between communication and computation.
Corollary 2.
Suppose Assumptions 24 hold. Let the stepsize satisfy the condition (9) given in Theorem 1, and denote by $\hat{K}$ the number of iterations required for Acc-FlexGT to reach an arbitrary accuracy $\epsilon>0$.
i) Strongly convex: Under the additional Assumption 1 with $\mu>0$, letting
(16) $\alpha=\lceil\frac{\max\left\{\ln 2,1/2\ln\left(n\beta L/\mu\right)\right\}}{\sqrt{1-\sqrt{\rho_{W}}}}\rceil,$
we have
(17) $\hat{K}=\mathcal{O}\left(\frac{L}{\mu}\log\frac{\left[V_{0}\right]}{\epsilon}+\frac{\sigma^{2}}{n\mu^{2}\beta\epsilon}\right).$
ii) Convex: Under the additional Assumption 1 with $\mu\geqslant 0$, letting
(18) $\alpha=\lceil\frac{\max\left\{\ln 2,1/2\ln\left(n\beta\right)\right\}}{\sqrt{1-\sqrt{\rho_{W}}}}\rceil,$
we have
(19) $\hat{K}=\mathcal{O}\left(\frac{\sqrt{L}}{\epsilon}+\frac{L\sigma^{2}}{n\beta\epsilon^{2}}\right)\mathbb{E}\left[V_{0}\right].$
iii) Nonconvex: Letting
(20) $\alpha=\lceil{\frac{\max\left\{\ln 2,1/2\ln\left(\beta\max\left\{n,R_{0}\right\}\right)\right\}}{\sqrt{1-\sqrt{\rho_{W}}}}}\rceil$
we have
(21) $\hat{K}=\mathcal{O}\left(\frac{L}{\epsilon}+\frac{L\sigma^{2}}{n\beta\epsilon^{2}}\right)\mathbb{E}\left[f\left(\bar{x}_{0}\right)-f^{*}\right].$
Proof.
The proof is similar to that of Corollary 1. By Lemma 1, we have $1-\bar{\rho}_{W}\geqslant 1/2$, and the topology effect part $\varUpsilon_{2}$ can be dominated when $\bar{\rho}_{W}$ is sufficiently small. Specifically, let $\hat{\gamma}$ denote the obtained upper bound of the stepsize in (9), and let the stepsize $\gamma$ satisfy
$\min\left\{\frac{\ln\left(\max\left\{2,\beta^{2}\mu^{2}\hat{K}/\varUpsilon_{1}\right\}\right)}{\beta\mu\hat{K}},\,\,\hat{\gamma}\right\}$
for the strongly convex case; and
$\min\left\{\frac{1}{\sqrt{\varUpsilon_{1}\hat{K}}},\,\,\hat{\gamma}\right\}$
for the convex and nonconvex cases to ensure that the centralized error part $\varUpsilon_{1}$ matches the linear or sublinear term. Then, we obtain the target complexity results in (17), (19), and (21).
Based on Corollary 2, we derive the communication and computation complexities of Acc-FlexGT, as shown in Table 3. Compared to FlexGT, Acc-FlexGT achieves a Pareto-optimal trade-off in terms of $\beta$ by selecting specific values of $\alpha$, as outlined in Corollary 2. In other words, there is no choice of $\beta$ that can simultaneously reduce both communication and computation complexity. This distinguishes our work from existing methods with accelerated communication [31, 10], which generally do not consider both communication and computational efficiency simultaneously, nor the trade-offs between them.
Remark 3.
Compared to existing results, Acc-FlexGT recovers DeTAG [31] as a special case when $\alpha=\beta$. Moreover, the iteration complexity (21) obtained for the nonconvex case is optimal in the sense that it matches the lower bound established in [31, 54], up to a logarithmic factor arising from the choice of $\beta=\alpha=\tilde{\mathcal{O}}\left(1/\sqrt{1-\sqrt{\rho_{W}}}\right)$. In the strongly convex case, our result improves the first term in (17) from $1/\sqrt{\epsilon}$ to $\log\left(1/\epsilon\right)$ compared to [10], as we use a constant stepsize instead of a stepsize that decays with iterations. To the best of our knowledge, the convex case is first considered in this work.
4 Convergence Analysis
This section presents a unified convergence analysis framework for the theoretical results stated in Theorem 1.
Proof sketch. The basic idea of the proof is illustrated in Fig. 3, which builds on the established contraction properties of the optimality gap for (strongly) convex objective functions (Lemma 2) and descent property for the nonconvex case (Lemma 3), both of which involve dependencies on the consensus and GT errors. Rather than using a direct recursive analysis, the key to bound these errors is to break the cyclic dependency between the consensus and GT errors (Lemmas 5 and 6) as highlighted by the arrows in the figure. For the (strongly) convex case, we achieve this by designing a suitable Lyapunov function in (8). Conversely, for the nonconvex case, we accomplish this through a straightforward method of accumulating the errors (Lemmas 7 and 8). Notably, the proof relies on a double-loop analysis. We first derive the upper bounds of the client divergence over the interval between two communications (inner loop) (cf., Lemma 4 in Appendix A). These bounds are subsequently utilized to establish the contraction properties of the consensus and GT errors during each communication round (outer loop). Following this proof sketch, we begin by introducing some key lemmas that are crucial for the proof of Theorem 1.
Figure 3: Basic idea of the unified convergence analysis framework. The solid arrows show the dependency of the terms. The dashed arrows indicate decoupling between the two accumulated terms. The dashed box outlines the Lyapunov and the accumulation methods for (strongly) convex and nonconvex objective functions, respectively.
4.1 Key Lemmas
For the (strongly) convex case, i.e., Assumption 1 hold, we have the following contraction property for the optimality gap.
Lemma 2 (Optimality gap).
Suppose Assumptions 14 hold. Let the step-size satisfy $\gamma\leqslant\frac{1}{4\beta L}$. Then, we have for all $k\geqslant 0$,
(22) $\mathbb{E}\left[\left\|\bar{x}_{\beta\left(k+1\right)}-x^{*}\right\|^{2}\right]$ $\leqslant\left(1-\frac{\mu\beta\gamma}{2}\right)\mathbb{E}\left[\left\|\bar{x}_{\beta k}-x^{*}\right\|^{2}\right]+\frac{\gamma^{2}\beta\sigma^{2}}{n}$
$\quad+\frac{3\gamma\beta L}{n}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]-\gamma\beta\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right].$
Proof.
According to the update rule of decision variable $\mathbf{x}$ (3a), we have
(23) $\mathbb{E}\left[\left\|\bar{x}_{\beta\left(k+1\right)}-x^{*}\right\|^{2}\right]$ $=\mathbb{E}\left[\left\|\bar{x}_{\beta k}-x^{*}\right\|^{2}\right]+\gamma^{2}\underset{=:S_{2}}{\underbrace{\mathbb{E}\left[\left\|\sum_{j=0}^{\beta-1}{\frac{\mathbf{1}^{\top}}{n}\mathbf{y}_{\beta k+j}}\right\|^{2}\right]}}$
$\quad-2\gamma\beta\underset{=:S_{3}}{\underbrace{\mathbb{E}\left[\left<\bar{x}_{\beta k}-x^{*},\frac{\mathbf{1}^{\top}}{n}\nabla F_{\beta k}\right>\right]}},$
where we used the fact that $\mathbb{E}\left[\nabla G_{\beta k}\right]=\nabla F_{\beta k}$.
Next, we bound the terms $S_{2}$ and $S_{3}$ in the above equation, respectively. For $S_{2}$, by (6) we have
(24) $\mathbb{E}\left[\left\|\sum_{j=0}^{\beta-1}{\frac{\mathbf{1}^{\top}}{n}\mathbf{y}_{\beta k+j}}\right\|^{2}\right]$ $\leqslant\frac{\beta\sigma^{2}}{n}+\mathbb{E}\left[\left\|\sum_{j=0}^{\beta-1}{\frac{1}{n}\sum_{i=1}^{n}{\nabla f_{i}\left(x_{i,\beta k}\right)}}\right\|^{2}\right]$
$\leqslant\frac{\beta\sigma^{2}}{n}+\frac{2\beta^{2}L^{2}}{n}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]+4\beta^{2}L\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right],$
where we have used Assumption 3 on bounded stochastic gradient variance and Assumption 2 on smoothness of the objectives. For $S_{3}$, using the convexity and smoothness of $f_{i}$ in Assumption 1 and 2, we have
[…]
$=\frac{1}{n}\sum_{i=1}^{n}{\mathbb{E}\left[\left<x_{i,\beta k}-x^{*},\nabla f_{i}\left(x_{i,\beta k}\right)\right>\right]}-\frac{1}{n}\sum_{i=1}^{n}{\mathbb{E}\left[\left<x_{i,\beta k}-\bar{x}_{\beta k},\nabla f_{i}\left(x_{i,\beta k}\right)\right>\right]}$
$\geqslant\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right]+\frac{\mu}{2n}\mathbb{E}\left[\left\|\mathbf{x}_{\beta k}-\mathbf{1}x^{*}\right\|^{2}\right]-\frac{L}{2n}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right].$
Substituting $S_{2}$ and $S_{3}$ in (23), and noticing that
(26) $-\mathbb{E}\left[\left\|\mathbf{x}_{\beta k}-\mathbf{1}x^{*}\right\|^{2}\right]$
$=-\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]-n\mathbb{E}\left[\left\|\bar{x}_{\beta k}-x^{*}\right\|^{2}\right]-2\sum_{i=1}^{n}{\left(\mathbb{E}\left[\left<x_{i,\beta k}-\bar{x}_{\beta k},\bar{x}_{\beta k}-x^{*}\right>\right]\right)}$
$\leqslant\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]-\frac{n}{2}\mathbb{E}\left[\left\|\bar{x}_{\beta k}-x^{*}\right\|^{2}\right],$
we get
(27) $\mathbb{E}\left[\left\|\bar{x}_{\beta\left(k+1\right)}-x^{*}\right\|^{2}\right]$
$\leqslant\left(1-\frac{\mu\beta\gamma}{2}\right)\mathbb{E}\left[\left\|\bar{x}_{\beta k}-x^{*}\right\|^{2}\right]+\frac{\gamma^{2}\beta\sigma^{2}}{n}$
$\quad+\left(\frac{2\gamma^{2}\beta^{2}L^{2}}{n}+\frac{2\gamma\beta L}{n}\right)\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]-\left(2\gamma\beta-4\gamma^{2}\beta^{2}L\right)\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right].$
Then, letting the stepsize satisfy $\gamma\leqslant\frac{1}{4\beta L}$, we complete the proof.
Next, we introduce lemmas for the nonconvex case, which play a key role in proving the convergence result in (12).
Lemma 3 (Descent lemma).
Suppose Assumptions 24 hold. Let the stepsize satisfy $\gamma\leqslant\frac{1}{4L\beta}$. Then, we have for all $k\geqslant 0$,
(28) $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]}$
$\leqslant\frac{4\left(\mathbb{E}\left[f\left(\bar{x}_{0}\right)\right]-\mathbb{E}\left[f^{*}\right]\right)}{\gamma\beta K}+\frac{4L^{2}}{n}\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}+\frac{2\gamma L}{n}\sigma^{2}.$
Proof.
By the smoothness of the objective function in Assumption 2, we get
(29) $f\left(\bar{x}_{\beta\left(k+1\right)}\right)$ $\leqslant f\left(\bar{x}_{\beta k}\right)+\left<\nabla f\left(\bar{x}_{\beta k}\right),\bar{x}_{\beta\left(k+1\right)}-\bar{x}_{\beta k}\right>+\frac{L}{2}\left\|\bar{x}_{\beta\left(k+1\right)}-\bar{x}_{\beta k}\right\|^{2}.$
Then, for the inner-product term on the right, by the recursion of decision variable $x$ in (3a) and noticing that $\mathbb{E}\left[\nabla G_{\beta k+j}\right]=\nabla F_{\beta k}$, we have
[…]
$\leqslant-\frac{\gamma}{2}\beta\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]+\frac{\gamma}{2}\mathbb{E}\left[\left\|\sum_{j=0}^{\beta-1}{\frac{\mathbf{1}^{\top}}{n}\left(\nabla F_{\beta k}-\nabla F\left(\mathbf{1}\bar{x}_{\beta k}\right)\right)}\right\|^{2}\right]$
$\leqslant-\frac{\gamma\beta}{2}\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]+\frac{\gamma\beta L^{2}}{2n}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right],$
Then, based on the upper bound of the GT error obtained in (24), we have
(31) $\mathbb{E}\left[f\left(\bar{x}_{\beta\left(k+1\right)}\right)\right]$ $\leqslant\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)\right]-\left(\frac{\gamma\beta}{2}-\gamma^{2}\beta^{2}L\right)\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]$
$\quad+\left(\frac{\gamma\beta L^{2}}{2n}+\frac{\gamma^{2}\beta^{2}L^{3}}{n}\right)\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]+\frac{\gamma^{2}\beta L}{2n}\sigma^{2}.$
Letting $\gamma\leqslant\frac{1}{8L\beta}$, we get
(32) $\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]\leqslant\frac{4\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(\bar{x}_{\beta\left(k+1\right)}\right)\right]}{\gamma\beta}+\frac{4L^{2}}{n}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]+\frac{2\gamma L}{n}\sigma^{2}.$
Summing the above inequality from 0 to k1k-1, we obtained the targeted results.
4.2 Proof of Theorem 1
With Lemmas 2 and 3 for optimality gap and objective function decrease, Lemmas 5 and 7 for consensus errors, and Lemmas 6 and 8 for tracking errors, we are now ready to prove Theorem 1.
Strongly convex case. Recalling the Lyapunov function (8) with
(33) $c_{x}=\frac{16\gamma\beta L}{n\left(1-\bar{\rho}_{W}\right)},\quad c_{y}=\frac{256\gamma^{3}\beta^{3}L\bar{\rho}_{W}}{n\left(1-\bar{\rho}_{W}\right)^{3}}.$
Then, with the help of Lemma 2, 5 and 6 and noticing that
$\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\leqslant 2L\left(f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right),$
we can get the contraction property for the Lyapunov function as follows:
(34) $\mathbb{E}\left[V_{\beta\left(k+1\right)}\right]$ $\leqslant\left(1-\max\left\{\frac{\mu\beta\gamma}{2},\frac{1-\bar{\rho}_{W}}{8}\right\}\right)\mathbb{E}\left[V_{\beta k}\right]$
$\quad+e_{1}\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right]+e_{2}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]+e_{3}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]$
$\quad+\frac{\gamma^{2}\beta\sigma^{2}}{n}+c_{x}\frac{8n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\sigma^{2}+c_{y}6n\sigma^{2},$
where
$e_{1}$ $=\frac{192n\gamma^{2}\beta^{2}L^{3}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}c_{y}-\gamma\beta,$
$e_{2}$ $=\frac{3\gamma\beta L}{n}+\frac{18\bar{\rho}_{W}L^{2}}{1-\bar{\rho}_{W}}c_{y}-\frac{3\left(1-\bar{\rho}_{W}\right)}{8}c_{x},$
$e_{3}$ $=\frac{4\gamma^{2}\beta^{2}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}c_{x}-\frac{1-\bar{\rho}_{W}}{4}c_{y}.$
Letting $e_{1}\leqslant-\gamma\beta/2$ and $e_{2},e_{3}\leqslant 0$ with the stepsize $\gamma$ satisfying
(35) $\gamma\leqslant\min\left\{\frac{1}{4\sqrt{2}\beta L},\frac{1-\bar{\rho}_{W}}{18\beta L\sqrt{\bar{\rho}_{W}}},\frac{\left(1-\bar{\rho}_{W}\right)^{2}}{40\beta L\bar{\rho}_{W}}\right\},$
we obtain the convergence result for the strongly convex case as shown in (10).
Convex case. Under Assumption 1 with $\mu=0$ and $x^{*}$ as one of the optimal solutions. Noticing that $e_{1}\leqslant-\gamma\beta/2$, we can rewrite the contraction property of the optimality gap in Lemma 2 as follows:
(36) $\mathbb{E}\left[\left\|\bar{x}_{\beta\left(k+1\right)}-x^{*}\right\|^{2}\right]$ $\leqslant\mathbb{E}\left[\left\|\bar{x}_{\beta k}-x^{*}\right\|^{2}\right]+\frac{\gamma^{2}\beta\sigma^{2}}{n}$
$\quad+\frac{3\gamma\beta L}{n}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]-\gamma\beta\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f^{*}\left(x^{*}\right)\right].$
Following the approach used in the strongly convex case, we apply the contracted Lyapunov function with the same values of $c_{x}$ and $c_{y}$ as in (33). By applying Lemmas 5 and 6 and ensuring that the stepsize satisfy (35), we have
(37) $\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right]\leqslant\frac{2\left(\mathbb{E}\left[V_{\beta k}\right]-\mathbb{E}\left[V_{\beta\left(k+1\right)}\right]\right)}{\gamma\beta}+\frac{2\gamma\sigma^{2}}{n}+\frac{3328\gamma^{2}\beta^{2}L\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}.$
Summing the above inequality from $k=0$ to K1K-1 and rearranging the terms, we obtain the result for the convex case as given in (11).
Nonconvex case. By combining the obtained Lemma 7 and Lemma 8 regarding the accumulated consensus error and GT error, respectively, we can further obtain
(38) $\left(1-\frac{768\gamma^{2}\beta^{2}L^{2}\bar{\rho}_{W}^{2}}{\left(1-\bar{\rho}_{W}\right)^{4}}\right)\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$
$\leqslant\frac{2\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)K}+\frac{32\gamma^{2}\beta^{2}\bar{\rho}_{W}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)^{3}K}+\frac{16n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{2}}\sigma^{2}+\frac{192n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}$
$\quad+\frac{3072n\gamma^{4}\beta^{4}L^{2}\bar{\rho}_{W}^{2}}{\left(1-\bar{\rho}_{W}\right)^{4}}\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]}.$
Letting the stepsize satisfy
(39) $\gamma\leqslant\min\left\{\frac{1}{4\beta L},\frac{1-\bar{\rho}_{W}}{14\beta L\sqrt{\bar{\rho}_{W}}},\frac{\left(1-\bar{\rho}_{W}\right)^{2}}{40\beta L\bar{\rho}_{W}}\right\},$
we get for all $K\geqslant 1$,
(40) $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$ $\leqslant\frac{C_{0}}{\left(1-\bar{\rho}_{W}\right)K}+\frac{416n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}$
$\quad+4n\gamma^{2}\beta^{2}\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]},$
where
(41) $C_{0}:=4\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{0}\right\|^{2}\right]+\frac{64\gamma^{2}\beta^{2}\bar{\rho}_{W}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)^{2}}.$
Then, by incorporating (40) into (28) in Lemma 3 and setting the initial consensus error in (41) to zero, we complete the proof.
5 Numerical Experiments
In this section, we report several numerical experiments to validate the theoretical findings of FlexGT and Acc-FlexGT on solving the distributed stochastic optimization problem (1), using both synthetic and real-world datasets. We compare FlexGT and Acc-FlexGT with several baseline methods, including DSGD [25], DSGT [37], DFL [29], K-GT [30], and SS-DSGT [10].
5.1 Synthetic Example
We consider the following distributed ridge regression problem over a network of $n=20$ nodes:
(42) $\underset{x\in\mathbb{R}^{p}}{\min}f\left(x\right)=\frac{1}{n}\sum_{i=1}^{n}{\underset{=:f_{i}}{\underbrace{\left(\mathbb{E}_{v_{i}}\left[\left(h_{i}^{T}x-\nu_{i}\right)^{2}+\frac{\mu}{2}\left\|x\right\|^{2}\right]\right)}}},$
where $\mu>0$ is the regularization parameter, $h_{i}\in\left[0,1\right]^{p}$ denotes the feature parameters of node $i$ with dimension $p=10$, and $v_{i}\sim\mathcal{N}\left(\bar{v}_{i},\sigma^{2}\right)$ with $\bar{v}_{i}\in\left[0,1\right]$. Therefore, the algorithm can obtain an unbiased noisy gradient $\nabla f_{i}\left(x_{i,t}\right)+\delta_{i,t}$ with $\delta_{i,t}\sim\mathcal{N}\left(0,\sigma^{2}\right)$ at each iteration $t$. Moreover, we set the stepsize according to the choice of $\alpha$ and $\beta$ as suggested by Theorem 1, i.e., $\gamma=c\left(1-\bar{\rho}_{W}\right)^{2}/\left(\bar{\rho}_{W}\beta L\right)$, where $c=10$ is a constant and the Lipschitz constant is set to $L=1$. We note that there is heterogeneity among the objectives $f_{i}$ of nodes due to the differences in $\{h_{i}\}_{i=1}^{n}$.
In Fig. 4, we present the communication and computation complexity of FlexGT to achieve an accuracy of $\epsilon=10^{-3}$ with $\alpha=1,2,\dots,8$ and $\beta=1,2,\dots,8$. The experiment was run 100 times for each parameter setting to obtain an average complexity. The results demonstrate that, for a fixed number of communication steps $\alpha$, increasing the number of computation steps per round $\beta$ reduces the communication complexity but increases the computation complexity (moving towards the upper left). However, when $\beta$ is fixed and $\alpha$ is increased, the computation complexity decreases initially, while the communication complexity first decreases and then increases. It is worth noting that the Pareto-optimal solutions (bold red lines) are achieved according to the value of $\alpha$ suggested by Corollary 2. For smaller $\mu$, more communication and computation are required, as shown in the plot on the right. These observations validate our theoretical findings and illustrate the trade-off between communication and computation in our distributed optimization algorithms.
(a) $\mu=0.001$
(b) $\mu=0.0001$
Figure 4: Communication and computation complexity of FlexGT algorithm to achieve $\epsilon=10^{-3}$ accuracy with $\alpha=1,2,\dots,8$ and $\beta=1,2,\dots,8$ on synthetic data. Each node is in an exponential graph of $n=20$ nodes with 5 neighbors.
In Fig. 5, we compare the convergence of the residual $\left|x_{k}-x^{*}\right|^{2}/\left|x_{0}-x^{*}\right|^{2}$ for the problem (42) between FlexGT, Acc-FlexGT, and baseline methods. Following the recommendation of Acc-FlexGT, we set the number of communication and computation steps according to (16) with $\alpha=\beta$. The results show that both FlexGT and Acc-FlexGT achieve superior convergence rates compared to other algorithms. This advantage is particularly evident in the scenario with weaker topological connectivity, i.e., when $\rho_{W}$ is larger, as shown in the plot on the right.
(a) $\left|\mathcal{N}_{i}\right|=5,\rho_{W}=0.74$
(b) $\left|\mathcal{N}_{i}\right|=3,\rho_{W}=0.88$
Figure 5: Comparison of convergence performance on synthetic data. Each node in an exponential graph of $n=20$ nodes with $\left|\mathcal{N}_{i}\right|$ neighbors.
5.2 MNIST Dataset
We conduct distributed training of three-layer neural networks with smooth activation functions for 10-class classification on the MNIST dataset [9]. The original training set is balanced, containing an equal number of samples for each label from 0 to 9. To introduce heterogeneous label distributions, we partition the training data across 8 GPUs connected over a ring graph, where each GPU receives samples corresponding to only 8 or 5 classes, as illustrated in the left column of Fig. 6. The average Kullback-Leibler divergence [22] for these two label distributions is 0.23 and 0.78, respectively, where a higher value indicates greater heterogeneity between each node’s local dataset and the overall training set.
Specifically, we evaluate the trained model on the full test dataset and compare the testing accuracy of FlexGT and Acc-FlexGT against baseline methods, including DSGD, DSGT, K-GT, and DFL. As shown in Fig. 6, with $\alpha=\beta=3$, when the data heterogeneity is high (top plots), FlexGT and Acc-FlexGT outperform the baseline methods in both communication and computation efficiency, with Acc-FlexGT achieving further performance improvements. Notably, the computation complexity of K-GT is worse due to simply skipping communications. These results illustrate the effectiveness of the proposed algorithms. Additionally, when data heterogeneity is relatively low (bottom plots), the gap between DSGD-based and GT-based algorithms becomes smaller, highlighting the robustness of GT in handling data heterogeneity.
Figure 6: Comparison of test accuracy against computation and communication steps on the MNIST dataset with heterogeneous label distributions. The top plots represent each node having data from 5 classes (higher heterogeneity), while the bottom plots correspond to each node having data from 8 classes (lower heterogeneity).
6 Conclusions
In this paper, we proposed FlexGT, a flexible snapshot GT method, along with its accelerated variant, Acc-FlexGT, for solving distributed stochastic optimization problems in non-i.i.d. scenarios. Leveraging a unified convergence analysis framework for strongly convex, convex, and nonconvex objective functions, we prove that FlexGT and Acc-FlexGT achieve linear or sublinear convergence rates with the best-known communication and computation complexity in most settings. In particular, Acc-FlexGT attains optimal iteration complexity in the nonconvex case (up to a logarithmic factor) and improves existing results in the strongly convex case. These findings highlight the inherent trade-off between communication and computation, with Acc-FlexGT achieving a Pareto-optimal solution when prior knowledge of the graph is available.
Future work includes exploring single-loop algorithms with randomized communication and computation strategies and evaluating the proposed methods in practical distributed optimization tasks, such as distributed training of large models and multi-robot system applications.
Appendix A Supporting Lemmas
This section presents several supporting lemmas for the proof of Theorem 1.
Lemma 4 (Bounding client divergence).
Suppose Assumptions 24 hold. Let the stepsize satisfy $\gamma\leqslant\frac{1}{4\beta L}$. Then, we get that for all $k\geqslant 0$
[…]
Proof.
By the $x$-update of FlexGT (3a), we have
(44) $\mathbb{E}\left[\left\|\mathbf{x}_{\beta\left(k+1\right)}-\mathbf{1}\bar{x}_{\beta k}\right\|^{2}\right]$ $\leqslant\mathbb{E}\left[\left\|\left(\bar{W}-\mathbf{J}\right)\tilde{\mathbf{x}}_{\beta k}-\gamma\bar{W}\sum_{j=0}^{\beta-1}{\mathbf{y}_{\beta k+j}}\right\|^{2}\right]$
$\leqslant 2\bar{\rho}_{W}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]+2\gamma^{2}\mathbb{E}\left[\left\|\sum_{j=0}^{\beta-1}{\mathbf{y}_{\beta k+j}}\right\|^{2}\right].$
Noticing that
(45) $\mathbb{E}\left[\left\|\mathbf{y}_{\beta k+t}\right\|^{2}\right]$ $\leqslant 2\mathbb{E}\left[\left\|\mathbf{y}_{\beta k}\right\|^{2}\right]+4n\sigma^{2}$
$\leqslant 4\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+4\mathbb{E}\left[\left\|\mathbf{1}\bar{y}_{\beta k}\right\|^{2}\right]+4n\sigma^{2}$
$\leqslant 4\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+8n\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]$
$\quad+8L^{2}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]+4\sigma^{2}+4n\sigma^{2}.$
Plugging the inequality above into (44) and letting the stepsize satisfy $\gamma\leqslant\frac{1}{4\beta L}$, we complete the proof.
Lemma 5 (Consensus error).
Suppose Assumptions 24 hold. Let the stepsize satisfy $\gamma\leqslant\frac{1}{4\beta L}$. Then, we have for all $k\geqslant 0$,
(46) $\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ $\leqslant\frac{1+\bar{\rho}_{W}}{2}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]+\frac{4\gamma^{2}\beta^{2}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+\frac{8n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\sigma^{2}.$
Proof.
According to the update rule of the decision variable in (3a), we have
(47) $\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ $\leqslant\frac{1+\bar{\rho}_{W}}{2}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]$
$\quad+\frac{\gamma^{2}\beta\left(1+\bar{\rho}_{W}\right)\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\sum_{t=0}^{\beta-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k+t}\right\|^{2}\right]},$
where we have used Young’s inequality [53]. Noticing that
(48) $\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k+t}\right\|^{2}\right]\leqslant 2\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+4n\sigma^{2},$
and letting the stepsize satisfy $\gamma\leqslant\frac{1}{4\beta L}$, we complete the proof.
Lemma 6 (Tracking error).
Suppose Assumptions 24 hold. Let the stepsize satisfy
(49) $\gamma\leqslant\min\left\{\frac{1}{4\beta L},\frac{1-\bar{\rho}_{W}}{14\beta L\sqrt{\bar{\rho}_{W}}}\right\}.$
Then, we have that for all $k\geqslant 0$,
(50) $\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ $\leqslant\frac{3+\bar{\rho}_{W}}{4}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+\frac{18\bar{\rho}_{W}L^{2}}{1-\bar{\rho}_{W}}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]$
$\quad+\frac{96n\gamma^{2}\beta^{2}L^{2}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]+6n\sigma^{2}.$
Proof.
Applying the recursion of FlexGT (3b), by Assumption 3, we have
(51) $\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ $\leqslant\frac{1+\bar{\rho}_{W}}{2}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+2n\bar{\rho}_{W}\sigma^{2}$
$\quad+\frac{\left(1+\bar{\rho}_{W}\right)\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\mathbb{E}\left[\left\|\nabla F_{\beta\left(k+1\right)}-\nabla F_{\beta k}\right\|^{2}\right]$
[…]
where we have used Young’s inequality and $\mathbb{E}\left[\nabla G_{\beta k}\right]=\nabla F_{\beta k}$. Then, noticing that
[…]
$=2\bar{\rho}_{W}\mathbb{E}\left[\left<\left(\bar{W}-\mathbf{J}\right)\nabla G_{\beta k},\nabla F_{\beta k}-\nabla G_{\beta k}\right>\right]\leqslant 2\bar{\rho}_{W}n\sigma^{2},$
and
$2\bar{\rho}_{W}\mathbb{E}\left[\left<\nabla F_{\beta\left(k+1\right)},\nabla F_{\beta k}-\nabla G_{\beta k}\right>\right]\leqslant\bar{\rho}_{W}\mathbb{E}\left[\left\|\nabla F_{\beta\left(k+1\right)}-\nabla F_{\beta k}\right\|^{2}\right]+\bar{\rho}_{W}n\sigma^{2},$
we get
(52) $\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta\left(k+1\right)}\right\|^{2}\right]$ $\leqslant\frac{1+\bar{\rho}_{W}}{2}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]+\frac{6\bar{\rho}_{W}L^{2}}{1-\bar{\rho}_{W}}\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]$
$\quad+\frac{6\bar{\rho}_{W}L^{2}}{1-\bar{\rho}_{W}}\mathbb{E}\left[\left\|\mathbf{x}_{\beta\left(k+1\right)}-\mathbf{1}\bar{x}_{\beta k}\right\|^{2}\right]+5n\bar{\rho}_{W}\sigma^{2},$
where we have used the smoothness of the objectives. Using Lemma 4 to bound the second last term and letting the stepsize satisfy (49), we complete the proof.
From the above lemma, it can be found that there is an interdependence between the consensus error and the GT error. To break the coupling between these two errors, we need the following lemmas. For convenience, we set the initial consensus error to zero, i.e., $x_{i,0}=x_{0}\in\mathbb{R}^{p},i\in\left[n\right]$.
Lemma 7 (Accumulated consensus error).
Suppose Assumptions 24 hold. Let the stepsize satisfy $\gamma\leqslant\frac{1}{4\beta L}$. Then, for all $K\geqslant 1$,
(53) $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$
$\leqslant\frac{2\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)K}+\frac{16n\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{2}}\sigma^{2}+\frac{8\gamma^{2}\beta^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{2}K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}.$
Proof.
Summing the obtained inequality in Lemma 5 from 0 to K1K-1, we get
(54) $\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta\left(k+1\right)}\right\|^{2}\right]}$
$\leqslant\frac{1+\bar{\rho}_{W}}{2}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}+\frac{8n\gamma^{2}\beta^{2}\bar{\rho}_{W}K}{1-\bar{\rho}_{W}}\sigma^{2}+\frac{4\gamma^{2}\beta^{2}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}.$
Adding $\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{0}\right\|^{2}\right]$ on both sides and noticing $\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}\leqslant\sum_{k=0}^{K}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$, we can obtain the target result, which completes the proof.
Lemma 8 (Accumulated tracking error).
Suppose Assumptions 24 hold. Let the stepsize satisfy (49). Then, for all $K\geqslant 1$,
(55) $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}$
$\leqslant\frac{4\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]}{\left(1-\bar{\rho}_{W}\right)K}+\frac{24n\sigma^{2}}{\left(1-\bar{\rho}_{W}\right)K}+\frac{96\bar{\rho}_{W}L^{2}}{\left(1-\bar{\rho}_{W}\right)^{2}}\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$
$\quad+\frac{384n\gamma^{2}\beta^{2}L^{2}\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{2}}\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]}.$
Proof.
Based on Lemma 6 and following a similar proof process as in Lemma 7, summing the result from 0 to K1K-1, we have
(56) $\sum_{k=0}^{K}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}$ $\leqslant\frac{3+\bar{\rho}_{W}}{4}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}+\frac{18\bar{\rho}_{W}L^{2}}{1-\bar{\rho}_{W}}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{x}}_{\beta k}\right\|^{2}\right]}$
$\quad+\frac{96n\gamma^{2}\beta^{2}L^{2}\bar{\rho}_{W}}{1-\bar{\rho}_{W}}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]}+6Kn\sigma^{2}.$
Adding $\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]$ on the both sides and noticing that
$\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]}\leqslant\sum_{k=0}^{K}{\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{\beta k}\right\|^{2}\right]},$
we can obtain the target result.